# Merlin: Multi-Stage Curriculum Alignment for Multilingual Encoder and LLM Fusion

### v2

@@ Abstract
Large language models excel in English but still struggle with complex reasoning in many low-resource languages (LRLs). Existing encoder-plus-decoder methods, such as LangBridge and MindMerger, raise the accuracy for mid and high-resource languages, yet large performance gap remains for LRLs. We present MERLIN, a two-stage model-stacking framework that applies a curriculum learning strategy—from general bilingual bitext to task-specific data—and adapts only a small set of DoRA weights. On the AfriMGSM benchmark MERLIN improves exact-match accuracy by +12.9 pp over MindMerger and outperforms GPT-4o-mini. It also yields consistent gains on MGSM and MSVAMP (+0.9 and +2.8 pp), demonstrating effectiveness across both low and high-resource settings.
Merlin: Multi-Stage Curriculum Alignment for Multilingual Encoder and LLM Fusion
@@ 1 Introduction
Large language models (LLMs) have achieved impressive results across a broad range of natural language processing (NLP) tasks (OpenAI et al., 2024; Jiang et al., 2024; Team et al., 2024a). Despite these advances, their performance often deteriorates significantly when dealing with low-resource languages (LRLs) (Adelani et al., 2024b; Uemura et al., 2024; Ojo et al., 2024). This shortfall is partly due to the limited online presence of LRLs and heavy focus on high or mid-resource languages during the pretraining of these models (Touvron et al., 2023; Aryabumi et al., 2024).
To improve reasoning capabilities for LRLs, the most straightforward approach is to adapt LLMs through continual pre-training (Alves et al., 2024; Ng et al., 2025; Buzaaba et al., 2025) or post-training (Ogundepo et al., 2025). However, these methods often depend on access to in-domain data, which poses a significant challenge for LRLs due to limited data availability. Additionally, fine-tuning LLMs—especially those with billions of parameters—incurs substantial computational costs, making these approaches less feasible in resource-constrained settings. (Dettmers et al., 2023; Han et al., 2024). LangBridge (Yoon et al., 2024) and MindMerger (Huang et al., 2024) address this challenge by leveraging rich multilingual representations from an external encoder as input to decoder-only LLMs. This design significantly enhances multilingual reasoning capabilities while substantially reducing both data and computational requirements. These approaches achieved comparable or even exceeded continual training baseline on the multilingual math reasoning dataset for the evaluated languages (Shi et al., 2022), including some mid-resourced languages (MRLs) such as Swahili [^1] We refer MRLs as languages with at least one billion monolingual corpus on the web (Kudugunta et al., 2023). Other approaches, such as QAlign (Zhu et al., 2024) and SLAM (Fan et al., 2025), focus on question alignment where an LLM is first trained to translate questions to English, followed by LoRA fine-tuning. However, our empirical experiments show that these approaches cannot effectively generalize to LRLs.
@@
In this paper, we introduce our Multilingual Embedding-Enhanced Reasoning for Language Integration Network called Merlin (Figure 1), a two-stage framework that integrates insights from model stacking methods with question alignment approaches. In the first stage, we learn a lightweight mapping layer to align a strong multilingual encoder with a frozen LLM through three successive translation tasks following the principles of curriculum learning from general domain bilingual data, followed by math domain bilingual data, and finally the task data. In the second stage, we specialize the LLM itself via parameter-efficient fine-tuning while keeping the learned mapping connector fixed. This design combines data-efficient supervised training with targeted adaptation of the LLM’s internal representations, yielding tighter encoder–decoder alignment without the cost of full-model retraining.
We evaluate MERLIN on various math reasoning datasets such as MGSM (Shi et al., 2022), MSVAMP (Chen et al., 2024), and AfriMGSM (Adelani et al., 2024b), an extension of MGSM to several low-resource African languages. MERLIN surpasses the strongest baseline, MindMerger, SLAM and QAlign, by more than +11.0 accuracy points on AfriMGSM and delivers consistent gains on MGSM with +2.8 performance gain to MindMerger, demonstrating its effectiveness for cross-lingual mathematical reasoning in multilingual settings, including LRLs. To assess generalization beyond mathematical reasoning, we evaluate MERLIN on natural language inference tasks using AfriXNLI (Adelani et al., 2024b), with +6.8 improvement in performance over MindMerger. To encourage further researches that emphasize the coverage of LRLs, we will release our code and models upon paper acceptance.
Model Bn Th Sw Ja Zh De Fr Ru Es En Mrl. Hrl. Avg
@@
Key Multilingual Approaches–Metamath 7B & mT5-XL
@@
Key Multilingual Approaches–Gemma 2 9B & NLLB-600M-distilled
@@
Key Multilingual Approaches–Metamath 7B & mT5-XL
@@
Key Multilingual Approaches–Gemma 2 9B & NLLB-600M-distilled
@@
Table 1: Performance on MGSM and MSVAMP. Bn/Th/Sw are mid-resource (MRL); the other seven languages are high-resource (HRL). Sub-headings (e.g. “Metamath-7B & mT5-XL”) specify the decoder base model and the multilingual encoder. LangBridge, MindMerger, and MERLIN use both components; SLAM and QAlign adopt only the listed decoder. $\Delta$ rows show MERLIN’s absolute gain over MindMerger ($\mathord{+}x$ for improvements, -$\mathord{-}x$ for drops).
@@ 2 Methodology
Given a lightweight connector that learns to map multilingual encoder states into the English-centric large language model embedding space, MERLIN’s primary goal is to improve representational alignment between the encoder and the LLM within the model-stacking architecture, thereby enabling the LLM to more effectively leverage the rich multilingual features provided by the encoder. This is accomplished through a two-stage process. In the first stage, Model Stacking—multilingual parallel data is leveraged to enhance the mapping capabilities of the connector module, also referred to as the mapping layer. This component is responsible for projecting the encoder’s multilingual representations into the LLM’s embedding space. In the second stage, Task Specialization—the focus shifts to the LLM itself. Here, a rank-constrained adaptation (e.g., DoRA(Liu et al., 2024)) is applied, tuning a small set of parameters to enable the decoder to reason effectively over the enriched cross-lingual input, conditioned on both the encoder outputs and the Mapping Layer. This approach not only enhances the mapping capability of the connector module but also enables the LLM to adaptively utilize the rich multilingual representations during the task-solving process. The following sections detail the objectives and data used in each stage.
2.1 Stage I: Model stacking Stage
@@
An initial bridge is established with a generic bitext from $L_{\text{tgt}}$, a generic sentence in the target language, to English (Huang et al., 2024). The decoder receives only the projected encoder states
@@
(b) Question Alignment (Align.)
@@ 2.2 Stage II: Task Specialization
Stage II internalizes the cross-lingual signal inside the decoder; enhances the integration of multilingual representations within the decoder and plays a critical role in improving performance on complex reasoning tasks. The connector $\sigma^{\ast}$, the encoder and all pretrained LLM weights remain fixed; only rank-constrained DoRA (Liu et al., 2024) matrices $\theta$ inside the decoder are activated. For every English task example $(q,a^{\star})$ the input takes the augmented form $(X_{f},T)(q)$ defined above. The adaptation parameters are obtained through
@@
Table 2: AfriMGSM performance across various LLMs and key multilingual approaches. We report Exact Match scores averaged over the displayed languages. We use the NLLB-600M distilled model as the encoder, and Gemma2-9b-it for MindMerger, MERLIN, and as the base model for QAlign. We use AfriTeVa-v2-Large as the encoder of LangBridge. Note that the released LangBridge source code does not accept NLLB architecture.
@@ 3.1 LLMs evaluated
We evaluate two closed models(GPT-4o-mini and GPT-4o)and four open LLMs, including Aya-101 (Üstün et al., 2024), MetaMath 7B, MetaMath-Mistral 7B (Yu et al., 2023), and Gemma 2 9B (Team et al., 2024b). Aya-101 is only evaluated for the NLI as it is one of the strongest open model baseline for low-resource languages.
3.2 Method baselines
LangBridge
(Yoon et al., 2024) injects a frozen multilingual encoder in front of a decoder-only LLM. The encoder’s last hidden states replace the LLM’s token embeddings after a single linear projection ($\approx$3 M parameters) learned on 100k generic English task-specific instruction tuning dataset; all LLM weights stay frozen, so inference cost matches the base LLM plus the encoder.
QAlign-LoRA
(Zhu et al., 2024) follows a two-phase recipe. Phase 1 trains the LLM to translate machine-translated questions into English, aligning cross-lingual prompts without changing answers. Phase 2 freezes the backbone and attaches LoRA adapters (<1 % parameters) that are fine-tuned on the original English task corpus.
SLAM
(Fan et al., 2025) profiles layer-wise activation patterns, selects the six bottom feed-forward blocks as language-sensitive, and fine-tunes only those ($\approx$7 % of parameters) on mixed English and translated data in a single stage; all higher layers and attention weights remain fixed.
MindMerger
(Huang et al., 2024) augments a decoder-only LLM with a frozen multilingual encoder connected through a two-layer MLP mapping head ($\approx$ 9 M parameters). Step 1 trains the head on generic translation pairs where answers are replaced by English references; Step 2 re-trains the same head on translated task question–answer pairs while the LLM and encoder stay frozen.
For the encoder component in multilingual-integration approaches, we couple both MindMerger and MERLIN with the distilled 600M-parameter NLLB encoder (Team et al., 2022). For LangBridge, we follow its original setup: AfriTeva-v2-Large (Oladipo et al., 2023) is used on AfriMGSM and AfriXNLI, whereas mT5-XL is employed for MGSM, because the publicly released LangBridge implementation only supports encoders that subclass the XXXEncoderModel.
@@ 3.3 Languages and Tasks
Our evaluation spans four multilingual benchmarks. AfriMGSM (Adelani et al., 2024b) extends MGSM with grade-school mathematics problems translated into 16 African languages, providing fine-grained coverage of more low-resource settings. AfriXNLI (Adelani et al., 2024b) offers natural-language inference data in the same 16 languages by translating ten domain splits from the original XNLI. The canonical MGSM corpus (Shi et al., 2022) supplies math QA instances across eight mid-resource languages. Finally, MSVAMP (Chen et al., 2024) introduces multilingual variants of SVAMP with 1000 algebra-word problems in eleven languages, testing cross-lingual reasoning beyond GSM-style arithmetic.
@@ 3.4 Experimental settings
Sub-stage Ia (Mapping). For every language we sample 9 000 general-domain English–target sentence pairs from the NLLB training corpus (Team et al., 2022). These bitext pairs supply a coarse bilingual anchor for the connector.
Sub-stage Ib (Question Alignment). To align non-English prompts with English semantics, we translate 20 000 English questions with NLLB 200-3.3B and sample 3 000 translated pairs per language. For the mathematics track (AfriMGSM and MGSM) the source questions come from MetaMathQA (Yu et al., 2024). For the NLI track (AfriXNLI) we instead sample premise–hypothesis pairs from translated MultiNLI (Williams et al., 2018), concatenate them as the “question”, translate, and again retain 3 000 pairs.
Sub-stage Ic (Augmentation). For each target language we extract a distinct set of 3 000 question–answer pairs from the training split of the task corpus: translated MetaMathQA for the math benchmarks (AfriMGSM, MGSM, and MSVAMP) and translated MultiNLI for the NLI benchmark (AfriXNLI).
Task-Specialization Stage. We fine-tune the LLM (via PEFT) on a second sample of 3 000 training instances per language drawn from the same translated sources as above. Using a separate set of examples helps reduce the risk of overlap with Sub-stage Ic and ensures there is no data leakage into the evaluation sets.  [^2] The detailed training configuration is stated on Appendix
@@ 4 Experimental Results
We assess MERLIN on both math reasoning and NLI benchmarks, comparing it against open LLMs, recent multilingual integration methods, and the strongest available closed models. For math reasoning, we use two datasets: MSGM and MSVAMP. To assess performance on low-resource African languages across both tasks, we employ AfriMGSM and AfriXNLI. All results are reported as exact-match or accuracy; higher is better.
@@
Table 3 reports accuracy on sixteen African languages plus English and French on NLI task. MERLIN attains $\mathbf{55.6}\,\%$ average accuracy on the African subset, improving on MindMerger by +6.8+6.8 points and on the base Gemma–2 9B model by +15.3+15.3 points, and marginally surpassing the closed GPT-4o-mini ($54.2\,\%$). The gains are most pronounced for Amharic (+10.0), Ewe (+11.5), Hausa (+10.2), Kinyarwanda (+15.2) and Xhosa (+6.6), confirming the value of the cross-lingual mapping curriculum in truly low-resource settings. A moderate decline is observed on HRLs such as English (–2.4) and French (–18.8), as well as on Igbo and Lingala. Nevertheless, the overall improvement across eleven of the sixteen African languages indicates that MERLIN successfully redistributes modelling capacity toward languages that benefit the most from additional cross-lingual supervision, while maintaining competitive performance on those already well supported.
5 Analysis

### v3

# Merlin: Multi-Stage Curriculum Alignment for Multilingual Encoder-LLM Integration in Cross-Lingual Reasoning

@@ Abstract
Large language models (LLMs) excel in English but still struggle with complex reasoning in many low-resource languages (LRLs). Existing methods align LLMs with multilingual encoders, such as LangBridge and MindMerger, raising the accuracy for mid and high-resource languages, yet large performance gap remains for LRLs. We present Merlin, a model-stacking framework that iteratively refines in 2-stages based on a curriculum strategy (from general to specific where general is bilingual bitext and specific is task-specific data) and adapts only a small set of DoRA weights. On the AfriMGSM benchmark Merlin improves exact-match accuracy by +12.9 pp over MindMerger and outperforms GPT-4o-mini by 15.2 pp. It also yields consistent gains on MGSM and MSVAMP (+0.9 and +2.8 pp), demonstrating effectiveness across both low and high-resource settings.
Merlin: Multi-Stage Curriculum Alignment for Multilingual Encoder-LLM Integration in Cross-Lingual Reasoning
@@ 1 Introduction
Large language models (LLMs) have achieved impressive results across a broad range of natural language processing (NLP) tasks (openai2024gpt4technicalreport; jiang2024mixtralexperts; geminiteam2024gemini15unlockingmultimodal). Despite these advances, their performance often deteriorates significantly when dealing with low-resource languages (LRLs) (adelani2024irokobenchnewbenchmarkafrican; uemura-etal-2024-afriinstruct; ojo2024goodlargelanguagemodels). This shortfall is partly due to the limited online presence of LRLs both in data availability and NLP tools, and heavy focus on high or mid-resource languages during the pretraining of these models (touvron2023llama2openfoundation; ranathunga_neural_2023; aryabumi2024aya).
To improve reasoning capabilities for LRLs, the most straightforward approach is to adapt LLMs through continual pre-training (alves2024tower; ng2025sea; buzaaba2025lugha) or post-training (ogundepo2025improving). However, these methods often depend on access to in-domain data, which is scarce for LRLs due to the lack of annotators, speakers, and linguistic experts. Additionally, fine-tuning LLMs—especially those with billions of parameters—incurs substantial computational costs and energy usage, making these approaches less feasible in resource-constrained settings (dettmers2023qloraefficientfinetuningquantized; han2024parameterefficientfinetuninglargemodels). LangBridge (yoon-etal-2024-langbridge) and MindMerger (huang2024mindmergerefficientboostingllm) address this challenge by leveraging rich multilingual representations from an external encoder as input to decoder-only LLMs. This design significantly enhances multilingual reasoning capabilities while substantially reducing both data and computational requirements. On the other hand, curriculum learning—a training paradigm in which models are exposed to data of increasing difficulty—shown to improve both convergence speed and final performance (soviany_curriculum_2022) and have been applied in recent works to enhance LLMs reasoning ability (ma-etal-2025-problem).
@@
In this paper, we introduce Multilingual Embedding-Enhanced Reasoning for Language Integration Network (Figure 1), which is abbreviated to Merlin. It is a two-stage framework that integrates insights from model stacking with question alignment. This design combines data-efficient supervised training with targeted adaptation of the LLM’s internal representations, yielding tighter encoder–decoder alignment without the cost of full-model retraining. Inspired by the principles of curriculum learning, in the first stage, we learn a lightweight mapping layer to align a strong multilingual encoder with a frozen LLM through three successive tasks starting with general domain bilingual data, followed by math domain bilingual data, and finally the task data. In the second stage, we fine-tune the LLM with parameter-efficient methods while keeping the mapping layer fixed, enabling the LLM to internalize and better exploit the learned representations during reasoning. We evaluate Merlin on various math reasoning datasets such as Multilingual Grade School Math Benchmark (MGSM) (shi2022languagemodelsmultilingualchainofthought), Multilingual Semantic VAriant Math Problems (MSVAMP) (chen2024breakinglanguagebarriersmultilingual), and AfriMGSM (adelani2024irokobenchnewbenchmarkafrican), an extension of MGSM to several low-resource African languages. Merlin surpasses the strongest baseline, MindMerger, LayAlign, SLAM and QAlign, by more than +11.0 accuracy points on AfriMGSM and delivers consistent gains on MGSM with +2.8 performance gain to MindMerger, demonstrating its effectiveness for cross-lingual mathematical reasoning in multilingual settings, including LRLs. To assess generalization beyond mathematical reasoning, we evaluate Merlin on natural language inference tasks using AfriXNLI (adelani2024irokobenchnewbenchmarkafrican), with +1.6 improvement in performance over MindMerger.
Model Bn Th Sw Ja Zh De Fr Ru Es En MRL HRL Avg
@@
Key Multilingual Approaches–Metamath 7B & mT5‑XL
@@
Key Multilingual Approaches–Gemma 2 9B & NLLB‑600M‑distilled
@@
Key Multilingual Approaches–Metamath 7B & mT5‑XL
@@
Key Multilingual Approaches–Gemma 2 9B & NLLB‑600M‑distilled
@@
Table 1: Performance on MGSM and MSVAMP. Bn/Th/Sw are mid‑resource (MRL); the other seven languages are high‑resource (HRL). Sub‑headings (e.g. “Metamath‑7B & mT5‑XL”) specify the decoder base model and the multilingual encoder. LangBridge, LayAlign, MindMerger, and Merlin use both components; SLAM and QAlign adopt only the listed decoder. $\Delta$ rows show Merlin’s absolute gain over MindMerger ($\mathord{+}x$ for improvements, -$\mathord{-}x$ for drops).
@@ 2 Methodology
Given a frozen multilingual encoder and a lightweight connector into the LLM’s embedding space, Merlin improves cross-lingual reasoning through a two-stage process. The first stage learns to align encoder outputs with the LLM representation space, while the second introduces a small adaptation inside the decoder to internalize this alignment for task solving. The subsections below describe the objectives and data used in each stage.
2.1 Stage I: Model Stacking
@@
An initial bridge is established with a generic bitext from $L_{\text{tgt}}$, a generic sentence in the target language, to English (huang2024mindmergerefficientboostingllm). The decoder receives only the projected encoder states
@@
(b) Question Alignment (zhu2024questiontranslationtrainingbetter) (Align.)
@@ 2.2 Stage II: Task Specialization
Stage II internalizes the cross-lingual signal inside the decoder; enhances the integration of multilingual representations within the decoder and plays a critical role in improving performance on complex reasoning tasks. The connector $\sigma^{\ast}$, the encoder and all pretrained LLM weights remain fixed; only rank-constrained DoRA (liu2024doraweightdecomposedlowrankadaptation) matrices $\theta$ inside the decoder are activated. For every English task example $(q,a^{\star})$ the input takes the augmented form $(X_{f},T)(q)$ defined above. The adaptation parameters are obtained through
@@
Table 2: AfriMGSM performance across various LLMs and key multilingual approaches. We report Exact Match scores averaged over the displayed languages. We use the NLLB-600M distilled model as the encoder, and Gemma2-9b-it for LayAlign, MindMerger, Merlin, and as the base model for QAlign. We use AfriTeVa-v2-Large as the encoder of LangBridge. Note that the released LangBridge source code does not accept NLLB architecture.
@@ 3.1 LLMs evaluated
We evaluate two closed models(GPT-4o-mini and GPT-4o) and four open LLMs, including Aya-101 (ustun2024ayamodelinstructionfinetuned), MetaMath 7B, MetaMath-Mistral 7B (yu2023metamath), and Gemma 2 9B (gemmateam2024gemma2improvingopen). Aya-101 is only evaluated for the AfriXNLI as it is one of the strongest open model baseline for low-resource languages.
3.2 Baselines Method Compared
We evaluate our method against five strong baselines: three model merging approaches—LangBridge (yoon-etal-2024-langbridge), MindMerger (huang2024mindmergerefficientboostingllm), LayAlign (ruan2025layalignenhancingmultilingualreasoning) and two methods leveraging in-domain question translation data, QAlign-LoRA (zhu2024questiontranslationtrainingbetter) and SLAM (fan2025slamefficientmultilingualreasoning). LangBridge injects a frozen multilingual encoder before a decoder-only LLM, replacing token embeddings with linearly projected encoder states ($\approx$3M parameters) learned from 100k English instructions, and inference cost equals that of the LLM plus the multilingual encoder. QAlign-LoRA first aligns multilingual prompts via question translation and then fine-tunes lightweight LoRA adapters (<1% parameters) on English tasks. SLAM fine-tunes the six lowest feed-forward layers ($\approx$7% parameters) identified as language-sensitive on mixed English and translated data, freezing higher layers and attention weights. MindMerger connects a frozen multilingual encoder to the LLM through a two-layer MLP head ($\approx$9M parameters), trained first on translation pairs with English references, then on translated QA pairs while keeping all base weights fixed. LayAlign integrates encoder features into multiple LLM layers via a trainable layer-wise aligner, trained in two stages using translation and task data while keeping both encoder and LLM frozen.
For the encoder component in multilingual‐integration approaches, we couple both MindMerger and Merlin with the distilled 600M‐parameter NLLB encoder (nllbteam2022languageleftbehindscaling). For LangBridge, we follow its original setup: AfriTeva-v2-Large (oladipo-etal-2023-better) is used on AfriMGSM and AfriXNLI, whereas mT5-XL is employed for MGSM, because the publicly released LangBridge implementation only supports encoders that subclass the XXXEncoderModel.
@@ 3.3 Languages and Tasks
Our evaluation spans four multilingual benchmarks. AfriMGSM (adelani2024irokobenchnewbenchmarkafrican) extends MGSM with grade-school mathematics problems translated into 16 African languages, providing fine-grained coverage of more low-resource settings. AfriXNLI (adelani2024irokobenchnewbenchmarkafrican) offers natural-language inference data in the same 16 languages by translating ten domain splits from the original XNLI. The canonical MGSM corpus (shi2022languagemodelsmultilingualchainofthought) supplies math QA instances across eight mid-resource languages. Finally, MSVAMP (chen2024breakinglanguagebarriersmultilingual) introduces multilingual variants of SVAMP with 1000 algebra-word problems in eleven languages, testing cross-lingual reasoning beyond GSM-style arithmetic.
@@ 3.4 Experimental Settings
Sub-stage Ia (Mapping). For every language we sample 9 000 general-domain English–target sentence pairs from the NLLB training corpus (nllbteam2022languageleftbehindscaling). These bitext pairs supply a coarse bilingual anchor for the connector.
Sub-stage Ib (Question Alignment).
To align non-English prompts with English semantics, we translate 20 000 English questions with NLLB 200-3.3B and sample 3 000 translated pairs per language. For the mathematics track (AfriMGSM and MGSM) the source questions come from MetaMathQA (yu2024metamathbootstrapmathematicalquestions). For the NLI track (AfriXNLI) we instead sample premise–hypothesis pairs from translated MultiNLI (williams2018broadcoveragechallengecorpussentence), concatenate them as the “question”, translate, and again retain 3 000 pairs.
Sub-stage Ic (Augmentation).
For each target language we extract a distinct set of 3 000 question–answer pairs from the training split of the task corpus: translated MetaMathQA for the math benchmarks (AfriMGSM, MGSM, and MSVAMP) and translated MultiNLI for the NLI benchmark (AfriXNLI).
Task-Specialization Stage.
We fine-tune the LLM (via PEFT) on a second sample of 3 000 training instances per language drawn from the same translated sources as above. Using a separate set of examples helps reduce the risk of overlap with Sub-stage Ic and ensures there is no data leakage into the evaluation sets. [^1] The detailed training configuration is stated on Appendix
@@ 4 Experimental Results
We assessed Merlin on both math reasoning and NLI benchmarks, comparing it against open LLMs, recent multilingual integration methods, and the strongest available closed models. For math reasoning, we use two datasets: MSGM and MSVAMP. To assess performance on low-resource African languages across both tasks, we employ AfriMGSM and AfriXNLI. All results are reported as exact-match or accuracy; higher is better.
@@
Table 3 reports accuracy on sixteen African languages plus English and French on NLI task. Merlin attains $68.7\,\%$ average accuracy on the African subset, improving on MindMerger by +1.6+1.6 points and on the base Gemma–2 9B model by +28.4+28.4 points, and marginally surpassing the closed GPT‑4o ($64.3\,\%$). The gains are most pronounced for Amharic (+5.6), Oromo (+4.6), Hausa (+2.5), and Kinyarwanda (+2.0), confirming the value of the cross‑lingual mapping curriculum in truly low‑resource settings. The overall improvement across thirteen of the sixteen African languages indicates that Merlin successfully redistributes modelling capacity toward languages that benefit the most from additional cross‑lingual supervision, while maintaining competitive performance on those already well supported.
5 Analysis of Language alignment and Effects on multilingualism